OpenStudy (jchick):

Someone please help!! Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6.

OpenStudy (jchick):

a. y + 3 = -6(x + 2) b. y + 2 = 6(x + 3) c. y - 3 = -6(x - 2) d. y - 2 = 6(x - 3)